### v2

# Research Paper

Abstract
Over the last decade, the Australian Dark Energy (OzDES) collaboration has used Reverberation Mapping to measure the masses of high redshift supermassive black holes. Here we present the final review and analysis of this OzDES reverberation mapping campaign. These observations use years of photometric and spectroscopic observations of Active Galactic Nuclei (AGN) in the redshift range and bolometric luminosity range . Both photometry and spectra are observed in visible wavelengths, allowing for the physical scale of the AGN broad line region to be estimated from reverberations of the H, MgII and CIV emission lines. We successfully use reverberation mapping to constrain the masses of super-massive black holes, and combine with existing data to fit a power law to the lag-luminosity relation for the H and MgII lines with a scatter of dex, the tightest yet identified, fit specifically for consistency with high redshift AGN. We fit a similarly constrained relation for CIV, resolving a tension with the low luminosity literature AGN by accounting for selection effects arising from finite survey length. We also examine the impact of emission line width and luminosity (related to accretion rate) in reducing the scatter of these scaling relationships and find no significant improvement over the lag-only approach for any of the three lines. Using these relations, we further estimate the masses and accretion rates of AGN with single epoch methods. We also use these relations to estimate the relative sizes of the H, MgII and CIV emitting regions, and find evidence that the MgII emission may occur further out than H. In short, we provide a comprehensive benchmark of high redshift AGN reverberation mapping at the close of this most recent generation of surveys, including light curves, time-delays, and a set of significantly improved radius-luminosity relations for use with high-redshift populations.
keywords:
galaxies: active – galaxies: nuclei – quasars: emission lines – quasars: general – quasars: supermassive black holes.1 Introduction
Every massive galaxy is thought to host a super-massive black hole (SMBH) at its core, with these objects believed to play an important role in galaxy evolution (Kormendy_2013). Despite their ubiquity, these objects remain poorly understood, with the exact mechanisms of their formation and growth within their host galaxies remaining as open questions (e.g. Volonteri_2021). Measuring the mass of these SMBHs over cosmic time is then of great interest, as it offers us insight into the history and evolution of these important objects.
As dark and compact objects, our ability to observe SMBHs by electromagnetic means relies on observations of their environment, and that ability drops rapidly with distance. This obstacle would frustrate our attempts to observe them into the cosmic past were it not for Active Galactic Nuclei (AGN), extremely luminous SMBHs with bright accretion disks that can easily outshine their entire host galaxy. Under the standard model (Urry_1995), these AGN consist of a central black hole / accretion disk ‘engine’ and a complex surrounding structure, including the fast orbiting material in the broad line region (BLR). As the luminosity, kinematics and structure of the AGN are all powered by the driving gravitational force of the central black hole, many of the AGN’s physical properties offer means of inferring the mass of the black hole.
AGN in the local universe can have their masses constrained by the kinematics of their surrounding stars, either directly tracking orbits (e.g. Schodel_2002) or the bulk stellar velocity dispersion (i.e. the ‘ relation’ Ferrarese_2000; Gebhardt_2000), but such techniques become challenging at larger distances where entire galaxies appear as AGN-dominated point sources. The previous decade has also seen considerable progress in imaging two of the nearest SMBHs with interferometry at millimetre wavelengths (e.g. EHT_2019), and in using Very Large Telescope interferometry to potentially access finer angular resolutions (GravityPlus_2022). At higher redshifts, the limits on angular resolutions can also be surpassed with reverberation mapping (RM), a technique in which temporal coverage of variability can substitute for a lack of spatial information (Blandford_McKee_1982; Peterson_1993).
Conceptually, reverberation mapping is straightforward: stochastic variations in the brightness of the AGN’s central engine drive variations in the brightness of gas in the surrounding broad-line region, but do so with a time delay associated with light travel time between the central engine and the BLR. Where we can distinguish signals from different regions, we can characterise the scale of this time delay and so infer the physical scale of the AGN (see Figure 1 for a sketch of this process with simplified geometry). Knowledge of this geometric scale, coupled with a measure of the velocity and assumptions about kinematics, provides a means of constraining the mass of the SMBH. As such, reverberation mapping of AGN has become the primary means of constraining SMBH masses at any appreciable redshift, dominating over all other methods beyond (Cackett_2021).
For most of the history of reverberation mapping, data have been limited to studies of a hand-full of AGN at a time, predominantly at low redshift (e.g. Peterson_1999; CIV_Peterson_2005; Denney_2006), limited by the difficulty of continuous photometric and spectroscopic observations over the required time scales. This has changed with the advent of the first generation of ‘industrial scale surveys’, in which reverberation mapping is performed on hundreds of AGN at redshifts probing into the deep cosmological past. Over the last decade, surveys like the Australian Dark Energy Survey (OzDES; King_2015) and the Sloan Digital Sky Survey (SDSS; SDSS-Shen_2015) have made regular photometric and spectroscopic measurements of well over AGN in the redshift range , dramatically increasing the number of distant AGN observations with optical RM data to a statistically significant sample, and pushing our constraints on SMBH masses well into the high redshift domain.
Reverberation mapping is observationally expensive, requiring many epochs of observation. However, black hole masses can be derived with single epoch spectral observations through an empirically observed power-law relationship between luminosity and reverberation timescale (or lag) for each line, the so-called ‘ Relationship’. The parameters of these relationships are calibrated using RM and they are reasonably well defined at low redshift, but the paucity of measurements of more distant AGN means that there has been difficulty constraining relations for high redshift AGN. One of the aims of OzDES has been to fill that observational gap. Here we present the complete OzDES RM sample and also make a comprehensive collation and comparison of contemporary reverberation mapping surveys, and create a catalogue of the SMBH masses derived from reverberation mapping.
The paper is organised as follows. We discuss the physical principles of reverberation mapping in Section 2, including methods and limitations. In Section 3 we introduce the OzDES and external data sets used in this paper, and then in Section 4 we introduce the methodology used by OzDES, including lag measurement procedures and quality cuts, used to define this data set. We then present the full suite of recovered lags, SMBH masses and other results from the OzDES RM program in Section 5. In Section 6.1 we compare the relationship constraints of OzDES against those of SDSS and a wide range of other surveys for the H, MgII and CIV lines, identifying where they agree or are in tension in the plane, and use a combined data set to tightly constrain the slope, offset and scatter for all three lines for samples similar to OzDES and SDSS. In Section 6.2 we confirm an absence of systematic biases from emission-line width or accretion rate for these fits. In Section 7 we derive a bolometric relation and use the different lag predictions for H, MgII and CIV as a probe of the relative scales of their emission regions. We then apply the updated relationships in Section 8 to derive single-epoch SMBH masses for AGN, and finally discuss the results of this paper, examining possible implications about the evolution of SMBH populations via our single epoch mass estimates, in Section 9.
2 Technique and limitations
2.1 Reverberation Mapping
The core of AGN reverberation mapping is based on arguments of the virialised orbit of the broad line region about the black hole, with the SMBH mass being correlated with the estimated radius of the BLR, , and its velocity dispersion as in equation 1. The complexities of the BLR’s unresolved geometry are captured in the broad uncertainties of the dimensionless ‘virial factor’, :111Mejia_Restrepo_2018 finds that virial factor may be anti-correlated with line width. In this work, we treat as being independent.
Though conceptually simple, each term in the numerator of equation 1 is non-trivial to measure for any given source. One hurdle is that the kinematics at play are dependent on the shape and dynamics of the BLR, the nature of which is as yet unresolved (Pancoast_2014a; Pancoast_2014b). This uncertainty is quantified by calibrating the population average virial factor, from population analysis of nearby AGN, the masses of which are estimated through use of the relation (Woo_2015; Grier_2013b), or more recently through dynamical modelling via velocity resolved reverberation mapping (Villafana_2023_dynamical; Shen_2024_dynamical). The is only loosely constrained for any individual source, as there is an inherent scatter of dex about , and this dominates the uncertainty in BLR RM derived masses (Woo_2015; SDSS-Shen_2023).
Measurements of are made from the spectroscopic broadening of the reverberating emission lines, either by their full-width half-maximum (FWHM), or their dispersion (i.e. the second moment of the line profile about its peak), with different virial factors being tuned for each (see Section 3.3 for examples of RM papers using either technique).
High redshift AGN are beyond the distances at which we have been able to angularly resolve , and so BLR RM instead uses the timescale at which light propagates through the AGN geometry to measure this size by proxy (see Figure 1). Variations in broad band photometric luminosity are dominated by the light emanating directly from the accretion disk ‘engine’, with these variations being echoed in the driven response in the emission lines of the BLR after some delay, , typically of order days to months. The BLR spectrum exhibits atomic recombination lines that can be easily distinguished from the broad AGN continuum emission, allowing us to identify this echo through spectroscopic observations even where our photometric observations are dominated by light from the accretion disk. With sufficiently tight cadence of photometric and spectroscopic observations, the delay (or lag) between driving continuum and driven emission line response, can be observed, allowing temporal resolution of observations to substitute for the lack of spatial resolution:
To perform reverberation mapping, a source must provide a clearly identifiable reverberating emission line from the BLR. At cosmological scales, the emission from different emission lines redshift in and out of typical spectroscopic wavelength ranges, and so different atomic emission lines are visible at near, moderate, and high redshift ranges.
For the nearest AGN (), optical RM is performed with H at and H at (e.g. HBETA_Bentz_2014). Beyond measuring the mean lag, some studies have even gone as far as using high signal to noise (SNR) measurements of hydrogen recombination lines to constrain the BLR structure and kinematics in further detail. In this “velocity resolved reverberation mapping” (e.g. Denney_2009; Grier_2013a), the reverberating line is treated as being affected by varying degrees of Doppler broadening from different velocities at different radii within the BLR, resulting in different parts of the line profile reverberating at different times.
At higher redshifts (), optical RM is performed with the MgII line at (e.g. MgII_Metzroth_2006). optical RM masses derived from MgII are complicated by iron contamination from the adjacent FeII emission lines, which can distort estimates of the line width unless accounted for (discussed in detail in Section 4).
At high redshift (), optical RM is performed with the CIV line at . CIV lag measurements have also been made by observing this line in the observer-frame UV range for low redshift and low luminosity sources (e.g. CIV_Rosa_2015; CIV_Peterson_2005; MgII_Metzroth_2006).
The relationship (Kaspi_2000), an empirically observed correlation between the luminosity of an AGN and its physical scale as recovered from RM, is parameterised as a power-law expressed as:
where is the rest-frame lag and is the monochromatic luminosity of the AGN at a line-specific wavelength: typically for H, for MgII, and for CIV. Though the relationship for H is well constrained and found to have small scatter ( dex; HBETA_Bentz_2013), recent surveys have suggested this scatter may be higher (e.g. HBETA_Du_2016; HBETA_Du_2018). The historical relations for MgII and CIV are found to have higher scatter than H (MgII_Zajacek_2020; Kaspi_2007).
The strong dependence of AGN lags on luminosity has given rise to the technique of stacking (Fine_2012_stacking; Fine_2013_stacking; Li_2017_stacking), in which sources of poor signal to noise but similar luminosity have their lag measurements combined. Through stacking, lag measurements can be made up in aggregate over many sources that would otherwise give poorly constrained results, such as done for the OzDES data in OzDES-Malik_2024.
There have been attempts to explain the diversity of lags through more expressive relationships, reducing the intrinsic scatter by adding predictive variables or dividing the AGN population into sub-groups based on the spectral properties (e.g. Mejia_Restrepo_2018; Mejia_Restrepo_2016). There is, for example, some evidence that lags are lower for highly accreting sources, above some critical threshold. (HBETA_Du_2016; HBETA_Du_2018). This is described by the dimensionless accretion rate, , given by equation 4, based on the thin-disk model of Shakura_1973.
Here, is the cosine of the inclination of the AGN, taken to be as an average value for the quasar inclinations in which the BLR and accretion disk are both visible (Seyfert - 1 type AGN), while is the bolometric luminosity at in units of and is the RM-derived SMBH mass in units of solar masses. The Super-Eddington Accreting Mass Black Hole (SEAMBH) collaboration use as a benchmark for the separation between low accretion “sub-Eddington” sources and the highly accreting “super-Eddington” sources (Du_2015; HBETA_Du_2016).
2.2 The Aliasing Problem
Though the uncertainties in within a single source are dominated by the dex population variability in , lag-recovery presents significant issues with bias and contamination of mass estimates. Cosmological time-dilation means that distant AGN have longer timescales of variability (Lewis_2023), meaning that RM requires such sources to be observed over a baseline of multiple years to adequately capture variations in their light curves. For the ground-based observations of most distant RM-campaigns, such multi-year observations are necessarily impacted by a seasonal windowing function that imposes 6-month gaps in observations. These seasonal gaps give rise to the problem of ‘aliasing’ (see Figure 2), which can yield spurious lag recoveries at yearly gaps (180 days, 420 days, etc.; OzDES-Penton_2021; OzDES-Malik_2022).
The dangers posed by aliasing are twofold; firstly, it degrades our ability to detect true lags that fall within seasonal gaps, and more dangerously it can lead to false positives arising within these same gaps (OzDES-Malik_2022). A common approach is to adopt weighting methods in combination with stringent selection criteria to remove sources with spurious lags. Such approaches reduce contamination from false positives, but at the cost of a drastically reduced sample, with typical acceptance ratios being of order in the OzDES sample. OzDES selections are based on agreement between competing methods (e.g. JAVELIN (JAVELIN-Zu_2010) and PyCCF (PyCCF-Mouyan_2018); discussed in detail in Section 4), the emergence of a single well constrained lag, and arguments about the physical reasonability of the overall lag posterior distribution (e.g. whether negative lags are properly excluded). More involved approaches rely on characterising the false positive rate (FPR) through simulation of sources (e.g. OzDES-Yu_2023; OzDES-Penton_2025). 222Recently, a new lag recovery code has been presented in the form of LITMUS,(McDougall_2025_LITMUS) which identifies false positives in a fully Bayesian way, but this method has not been used for any lag recoveries used in this paper.
3 Data
3.1 OzDES
The OzDES project (OzDES-DR0-Yuan_2015; OZDES-DR1-Childress_2017; OZDES-DR2-Lidman_2020) was a parallel project of the Dark Energy Survey (DES), providing spectroscopic follow-up of sources imaged photometrically by DES. DES itself performed a 6-year imaging survey on the CTIO Blanco 4-metre Telescope in Chile. In addition to a wide-field survey, DES repeatedly imaged ten fields to search for supernovae (field coordinates given in Table 2 of OzDES-Spec-Smith2018). These supernova fields had weekly imaging (during each 6 month season) for 5 years from 2013-2018 and a sixth year with lower cadence in 2019 (for details see OZDES-DR2-Lidman_2020), using DECam in and bands (Flaugher_2015). Primarily, these fields were used to obtain supernova light curves, but they also provide an extensive data set for monitoring the variability of other sources, such as AGN. For ease of access we provide the photometry relevant for AGN reverberation mapping with this paper [link to be provided upon acceptance]. The raw images and catalogues from the whole DES survey are available at https://des.ncsa.illinois.edu/releases/dr2.
To complement the DES photometry, OzDES obtained monthly spectroscopy using the AAOmega spectrograph fed by the Two Degree Field (2dF) fibre positioner (Lewis_2002_2df) on the Anglo-Australian Telescope (AAT; for details see OZDES-DR2-Lidman_2020). The 2dF instrument has 400 optical fibres that can be positioned on targets within a 2-degree diameter field of view, and about a quarter of those were placed on AGN in each exposure. Using a spectral resolution of to and a wavelength range of to OzDES monitored a total of AGN over six years, gathering between and epochs on each AGN. OzDES typically observed for five months between August and January, perforce leaving large seasonal gaps in the time series data while the fields passed near to the sun. The spectra are publicly available from https://docs.datacentral.org.au/ozdes/overview/dr2/.
3.2 Light Curve & Spectral Calibration
For reliable reverberation mapping, a multi-stage calibration procedure of the photometric and spectroscopic measurement is necessary. In OzDES, we make use of the calibration procedures outlined in OzDES-Hoormann_2019, with the exception of line flux and width measurements for the MgII line for which we use the procedure of OzDES-Yu_2023. In this section we briefly review this calibration pipeline, but direct the reader to the original papers for more detail.
DECam measurements provide fluxes at an irregular but approximately weekly cadence in the filters, though in OzDES RM we make use of only those in the , and bands. In the OzDES-Hoormann_2019 pipeline, spectroscopic epochs are discarded if they are of low quality (require quality flag 4 or greater, see OzDES-DR0-Yuan_2015) or lack a co-temporal calibrating magnitude. Epochs are then discarded as outliers if they differ by more than magnitudes from the source’s mean in the same year and same band pass, with the remaining measurements being averaged if they occur on the same night. An additional filter-dependent calibration uncertainty is also added to the variance as per Burke_2017.
Each spectral epoch is calibrated, both converting from photon counts to flux values and correcting for systematic distortions of the spectrum, by comparing the spectra to the , and photometric measurements. Synthetic photometric measurements are made by integrating over the product of the spectrum and band-pass, and a ratio taken between synthetic and measured brightness. Describing this factor as a function of wavelength, a quadratic is fit across the three measured ratios, acting as a measure of the relative throughput. The entire spectrum was warped to remove this bias (e.g. Figure 3). After warping, spectra are co-added if they share the same night.
Once the spectrum is calibrated, line fluxes and widths are calculated by isolating a window of rest-frame wavelengths in which the reverberating line occurs. By linearly interpolating between the flux at the window boundaries, the local continuum is subtracted and the total flux, line dispersion and line full width half maximum estimated. Uncertainties in this value are found by Monte Carlo variation of the window boundaries and the flux measurements within bounds of uncertainty.
Following the approach of HBETA_Du_2016, we apply a correction to the monochromatic luminosities of all H sources by way of the empirical scaling relationship provided by Equation 1 in shen_2011_stellarcontamination. A more thorough approach would be to decompose the spectra into host galaxy starlight and AGN activity, but is not crucial in this paper as the OzDES H sample’s small size means that it offers only weak constraining power to the parameters (see Section 6.1.1).
For the MgII sources we add an additional step to account for the iron emission lines that flank and coincide with the MgII emission line, which, if not accounted for, can contaminate measurements of line width. In their analysis of the OzDES MgII lags, OzDES-Yu_2023 account for this by fitting iron spectral templates (Template_Tzuzuki_2006; Template_Salviander_2007; Template_Vestergaard_2001) in a Bayesian fashion using the python Monte Carlo Markov Chain (MCMC) package emcee. Spectral fits for the emission lines were visually inspected and poorly fitted epochs were removed before lag-fitting with JAVELIN.
Following the line width pipeline of OzDES-Hoormann_2019, the measurements presented here are made using the spectrum averaged over all epochs and use line dispersion as a measure of the line width. All fluxes are converted to luminosities assuming a CDM cosmology with parameters km s-1Mpc-1, and .
3.3 External data sets
In addition to presenting our recovered lags, we make use of a number of sources from the literature, both for comparison and to improve the constraining power of the relationship. A summary of these sources are listed in Table 1 The bulk of these lags are those recovered from the Sloan Digital Sky Survey (SDSS-Shen_2023), a contemporary to OzDES at a similar ‘industrial scale’ and redshift range. In Section 4.2.4, we describe how the SDSS lags use dissimilar selection criteria to OzDES, including the use of different software as their primary detection method. For the sake of homogeneity in this paper, we adopt the JAVELIN lags for SDSS and apply an analogue of our own selection criteria.
| H Sources | ||
| Source | No. AGN | Redshift Range |
| OzDES | ||
| SDSS | ||
| Bentz Collection | ||
| SEAMBH | ||
| LAMP | ||
| Misc | ||
| MgII Sources | ||
| OzDES | ||
| SDSS | ||
| Zajacek | ||
| CIV Sources | ||
| OzDES | ||
| SDSS | ||
| Kaspi | ||
For low redshift lags associated with the H line, we also supplement our data with the external lags utilised in the analysis of OzDES-Malik_2023. These include lags from: the Lick AGN Monitoring Project (LAMP) (HBETA_U_2022), a low redshift () study of statistically diverse AGN; the SEAMBH survey, which targets highly accreting high luminosity sources known to produce lower reverberation lags (HBETA_Du_2016; HBETA_Du_2018; HBETA_Hu_2021); and a collection of low redshift sources collated by HBETA_Bentz_2013, coupled with similar sources from HBETA_Fausnaugh_2017; HBETA_Bentz_2009; HBETA_Bentz_2013; HBETA_Bentz_2014; HBETA_Bentz_2016a; HBETA_Bentz_2016b and HBETA_Bentz_2023. Also included are a number of additional sources from HBETA_Zhang_2019; HBETA_Li_2021; HBETA_Lu_2016; HBETA_Pei_2014; HBETA_Rakshit_2019. Sources prior to 2019 and not associated with the SDSS collaboration are drawn from a convenient collation by HBETA_Martinez-Aldama_2019.
The combination of a larger sample size and good signal-to-noise yields a larger set of lags associated with the MgII line in the OzDES intermediate redshift sample than are recovered in the low redshift H sample (OzDES-Yu_2021; OzDES-Yu_2023). In total, OzDES MgII lags are available, which we supplement and contrast with the set of SDSS MgII lags, a smaller set of lags from MgII_Zajacek_2021 and a wider collection of literature sources presented by MgII_Zajacek_2020. This collection is comprised of lags collated from MgII_Lira_2018; MgII_Metzroth_2006; MgII_Czerny_2019 and earlier SDSS lags from MgII_Shen_2016; MgII_Shen_2019. We use the entire set of lags as listed by MgII_Zajacek_2020 when analysing the collection by itself, but when using these data in combination with other surveys we defer to the SDSS measurements when the same source has been reported twice.
For the CIV lags, we compare the OzDES results presented in OzDES-Penton_2025 with the latest SDSS CIV lags from SDSS-Shen_2023 and a diverse collection of earlier literature sources, which we label as as ‘The Kaspi Collection’. It is worth noting that the Kaspi Collection includes SDSS lags from SDSS-Grier_2017; CIV_Grier_2019; MgII_Shen_2019, as well as earlier OzDES lags from OzDES-Hoormann_2019. As we already analyse more recent and complete SDSS results independently, we do not include these sources in our analysis of the Kaspi Collection.
The sources of the Kaspi Collection can be broadly separated into two groups: high redshift, high luminosity AGN with the CIV line measured in the optical band by MgII_Lira_2018 and CIV_Kaspi_2021, and a number of nearby low luminosity sources from CIV_Peterson_2005, MgII_Metzroth_2006 and CIV_Rosa_2015 in which the CIV line is observed in the observer-frame UV. We separate these two groups and label them as ‘Kaspi High-Z’ and ‘Kaspi Low-Z’. As we discuss further in Section 6.1.3, these two sub-samples have different statistical properties, and treating them as a single group runs afoul of tensions when constraining the relationship.
When considering the velocity in equation 1, there are two decisions about how this is interpreted from the quasar spectra: whether to use the dispersion / variance of the line profile (disp) or its full-width half maximum (FWHM), and whether to make these measurements from the spectrum as time-averaged over all epochs (mean-spec) or by its root-mean-squared over all epochs (RMS-spec). Different papers in the literature occupy all four quadrants of this decision tree. Note that the disp/FWHM decision represents a different choice of physical measurement, and so leads to a correspondingly different virial factor, while the mean-spec/rms-spec decision is purely about which gives the clearer measurement.
The dispersion approach tends to be the preferred approach in more modern surveys (e.g. SDSS-Shen_2023; HBETA_Rakshit_2019; HBETA_Zhang_2019), while both mean-spec (e.g. HBETA_Li_2021) and RMS-spec (e.g. Lu_2019) approaches are still in active use. The RMS spectrum is preferred for high SNR sources to isolate the time-varying component of the signals (e.g. HBETA_Bentz_2014; HBETA_Pei_2014), but many surveys will measure line widths in all, or at least multiple, ways (e.g. SDSS-Shen_2023; HBETA_Fausnaugh_2017).
In this work, as with past OzDES RM papers, we use the dispersion-based definition of line width, as measured on the mean spectrum, as our measure of velocity rather than the FWHM, and default to this measurement for literature measurements where available in the source paper. In making the decision between mean-spec and RMS-spec, we defer to the preferred method of the source paper, e.g. mean spectrum for sources from HBETA_Bentz_2013, RMS spectra for sources from SDSS-Shen_2023 etc. For source-papers in which only the FWHM line widths are provided, we follow the lead of CIV_Kaspi_2021 and use based on an assumption of a Gaussian line profile.
4 Methods
4.1 Lag Recovery
The central element of reverberation mapping is identifying and constraining the delay between two light curves, a task that is complicated by the stochastic variability in the underlying signal, and a host of obstacles arising from the windowing function. These issues have led to a wide range of competing techniques and software for computing the lag in AGN-RM. These can be broadly sorted into two categories: the first containing non-parametric techniques that assume no underlying statistical behaviour and fit to only the light curves, and the second being those that model the AGN variability as a Gaussian process.
Modelling the AGN light curve as a Gaussian process is physically motivated, as they are known to show consistent patterns in their variability (MacLeod_2010; Zu_2013; Kozlowski_2016), but the associated fitting procedures are computationally expensive and fraught with numerical and statistical obstacles (OzDES-Penton_2021; OzDES-Malik_2022; Read_2019). As a result, non-parametric methods like PyCCF (PyCCF-Mouyan_2018) and PyROA (pyroa-ferus_2021) are still in use as more robust, if less precise, alternatives to full Bayesian modelling. The SDSS team uses PyROA for their ‘primary’ lag recovery method, while OzDES use a combination of the GP-based JAVELIN and the non-parametric PyCCF for validation and source selection cuts. For the interested reader, in Appendix A we provide an introduction to these fitting methods, their statistical basis, and numerical shortfalls.
4.2 Post-Recovery Quality Cuts
The stochastic nature of AGN light curves means that non-physical false positive lags are a possibility even in the presence of arbitrarily good measurements. Coupled with matters of measurement noise, calibration error and, in multi-year surveys such as OzDES, the impacts of aliasing from the seasonal windowing function, the contamination rate from such false positives can outstrip physically meaningful lag recoveries by near to an order of magnitude if unaccounted for.
Bespoke analysis can be used to assess reliability of a handful of sources at a time, but this approach is untenable in industrial-scale studies like OzDES. Instead, it is necessary to use a general, widely applicable criteria for selecting and post-processing the lag recoveries of the sample sources at large to remove spurious results (OzDES-Penton_2021; OzDES-Malik_2022).
The methods for suppressing the impacts of aliasing through quality cuts varies significantly between different studies. In this section we outline the quality cuts used in each of the previous OzDES RM papers as well their underlying motivation, and present also a set of selection criteria to apply to the SDSS RM sample of SDSS-Shen_2023 aimed at drawing from their data a set of sources with selection effects analogous to those of OzDES.
4.2.1 Quality Cuts for H Lags
OzDES tracked AGN sources in the redshift range for H line visibility, with only sources passing the quality cuts outlined in OzDES-Malik_2023. The low luminosities for these sources give lags in the day range, meaning only a small fraction of the brightest sources are expected to be heavily impacted by aliasing, which becomes more severe when the true lag is ‘off season’, i.e. at etc. (OzDES-Malik_2022). The OzDES quality cuts for H are that each source must have:
1.
Uncertainties from the JAVELIN recovered lag are less than ,
2.
Lags from PyCCF and JAVELIN agree to within of the JAVELIN uncertainties,
3.
The maximum correlation from PyCCF ( in equation 12) gives , and
4.
The PyCCF false positive rate (p-value) is .
The first cut selects only well constrained lags, the second leverages the robustness of the Interpolated Cross-Correlation Function (ICCF) to account for spuriously constrained JAVELIN false positives, and the last two cuts ensure that we keep only lags that do not favour the null hypothesis of there being no relation between the two light curves.
4.2.2 Quality Cuts for MgII Lags
The majority of OzDES AGN sources are observed in the redshift range that allow for visibility of the MgII line, with of its target AGN being in the redshift range where the MgII line is visible in range of OzDES data. These MgII sources exist over a wide redshift range, and the range of observed lags significantly overlaps with the year ‘danger-zones’ for aliasing effects. These high redshift AGN lags require strict and principled quality cuts to avoid contamination of the relationship. To this end, OzDES-Yu_2023 applies two distinct sets of selection criteria, one based on the posterior distribution weighting approach of CIV_Grier_2019, and another based on the simulation based false positive rate estimation of OzDES-Penton_2021.
In the SDSS approach, the marginalised lag posterior distribution , as recovered by PyROA, JAVELIN or PyCCF, is attenuated by a weighting function that down-weights lags that correspond to the poorly constrained seasonal gaps. This weighting function is the convolution of two components. The first term, based on the fraction of observations that overlap between the two light curves after shifting by some lag, down-weights ‘off-season’ lags in seasonal gaps where this overlap is small. The second term is the auto-correlation function of the continuum signal (ACFCont), which smooths and widens the ‘on-season’ peaks of the weighting function to account for the continuous nature of the light curve. The entire posterior is then convolved with Gaussian smoothing kernel of width days (equation 5):
where represents the convolution operator. The square bracketed terms represent the weighting function while the convolved normal distribution represents the final smoothing.
This ad-hoc approach suppresses lags that are in danger of being the result of aliasing, retaining only those that give an extremely high likelihood. Different quality cuts have been used in past analysis based on this weighting, including discarding sources in which the weighting reduces the posterior evidence by too large of a fraction (e.g. CIV_Grier_2019).
In OzDES-Yu_2023, the OzDES analysis reports lags from the un-weighted distribution as produced by JAVELIN, but uses the weighted distribution to identify the ‘primary’ peak as being the highest likelihood mode in the weighted-distribution . Sources are retained only if the unweighted distribution has:
1.
More than % of its evidence (posterior distribution integral) contained within the primary peak,
2.
The width of this peak, as measured between the 16th and 84th percentiles, be less than days, and
3.
This peak be in agreement with the unweighted PyCCF distribution to within .
OzDES-Yu_2023 also includes quality cuts modelled after the more rigorous simulation-work of OzDES-Penton_2021, in which simulations of mock DRW signals are used to characterise false positive rates, and quality cuts tuned to remove suspicious sources. These cuts, designed specifically for unweighted JAVELIN posterior distributions and OzDES-like observations, require:
1.
That the standard deviation of the JAVELIN posterior distribution be less than days,
2.
That the separation of the JAVELIN median lag & peak likelihood lag be within days of one another, and
3.
That the lags from PyCCF and JAVELIN agree to within days.
Of the available MgII sources, pass both sets of quality cuts and produce reliable lags, of the initial sample, with an estimated false-positive rate in this post-cut sample of (i.e. false positive out of the ).
4.2.3 Quality Cuts for CIV Lags
OzDES tracked high redshift AGN with CIV lines visible for reverberation mapping. These distant sources represent the earliest and most luminous AGN in the OzDES RM sample, and, owing to the anti-correlation of AGN luminosity and optical variability (MacLeod_2010), the sources with the weakest AGN variability. These sources are also observed with the shortest rest-frame time window due to higher time dilation at these higher redshifts. In combination, lag recovery for the CIV sample is a difficult process with an outsized risk of false positives if approached naively. For CIV RM of the OzDES sample, OzDES-Penton_2021 provide a series of quality cuts to maximise the reliability of constraints and minimise the rate of false positives, tuned on mock simulations of OzDES-like data. Under these cuts, sources are sorted into quality levels of ‘bronze’, ‘silver’, ‘gold’, or are rejected completely based on the properties of their marginalised lag posterior distribution and the lag recovery results from the ICCF method (examples of gold and bronze CIV sources can be seen in Figure LABEL:fig:_CIV_example_curves). These cuts are:
1.
That the JAVELIN posterior rule out to within ,
2.
That the lag of JAVELIN and ICCF lags agree to within ,
3.
That the JAVELIN lag posterior median and maximum posterior value be similar (within for bronze quality, for silver and for gold),
4.
That the JAVELIN posterior is strongly constrained to a single peak, with a high fraction ( and for bronze, silver and gold, respectively) of posterior density / MCMC chain samples falling within a single mode.
After these cuts, OzDES-Penton_2025 recovered CIV lags, with each at the gold and silver confidence level and at bronze. This analysis included a re-examination of the two CIV lags from OzDES-Hoormann_2019 with more years of observations.
4.2.4 SDSS Sub-Sampling Criteria
In the final data release for the Sloan Digital Sky Survey’s 7-yr RM campaign, SDSS-Shen_2023 constrain and select lags in a fundamentally different way to OzDES. Rather than using simulation-based estimates of the false positive rate, SDSS instead allow each source to fit for both positive and negative lags, and then apply the aliasing mitigation weighting strategy outlined in Section 4.2.2. Under the rationale that any lag recovered at must necessarily be a false positive, they use this as an estimate of the population level false positive rate, using this to tune their selection criteria. To this end, they make use of only one quality cut: that the ICCF value (equation 12) must be as evaluated at the lag recovered by their chosen primary lag recovery method PyROA. This is similar to the selection criteria of the OzDES H analysis of OzDES-Malik_2023, which requires at the lag as recovered by JAVELIN.
By contrast to SDSS, OzDES treats this lag search range as a true prior, examining only the physically reasonable positive lag domain and applying stringent cuts to minimise the effect of aliasing on each individual source. This produces a more restricted sample with fundamentally different statistical properties, which can be seen in the tension between the SDSS and OzDES constraints. To make joint fits to the relation using both the SDSS and OzDES data sets, we required a set of lags with relatively homogeneous selection properties. To that end, we derive an alternate set of lags from the published SDSS data of SDSS-Shen_2023 that differ from their sample in two ways: firstly, in that we use the lags as recovered from JAVELIN instead of their choice of PyROA, and secondly that we apply a number of quality cuts to this alternate sample to bring it more in line with the rationale of the OzDES selections.
Our sub-sampling criteria for the SDSS sources are:
1.
That recovered lags must be positive, at .
2.
That results should be robust across methods, such that JAVELIN, ICCF and PyROA all agree in peak lag to within of the JAVELIN uncertainty,
3.
That the fit should be statistically significant, as measured by setting evaluated at the JAVELIN modal lag,
4.
That the anti-aliasing down-weighting weighting scheme must not overly distort the lag probability distribution. We require that the posterior evidence rejected (fraction of MCMC chain samples in their terminology) by the SDSS anti-aliasing weighting must be .
For step 2, in cases where the JAVELIN lag peak is strongly constrained, we replace the JAVELIN confidence interval in this test with days for H or days for CIV and MgII, following a similar reasoning to the cuts of OzDES-Penton_2021. After applying these cuts, we retain a sub-sample with H lags, MgII lags and CIV lags from the SDSS sample. JAVELIN’s MCMC sampler is known to exhibit strong aliasing artefacts in seasonal signals (see McDougall_2025_LITMUS), meaning that its use by itself can give misleading results. However, JAVELIN’s use in concert with specifically tuned quality cuts produces a set of reliable lags with a low FPR, as validated by the simulations of OzDES-Penton_2021. We use SDSS’s JAVELIN lags rather than their PyROA lags, as the OzDES quality cuts are tuned specifically for JAVELIN’s systematics.
The motivation of this sub-sampling is to sacrifice much of the SDSS sample’s statistical power by cutting many sources, including discarding what is likely a large fraction of true positives, to yield a smaller sample that has OzDES’s high reliability on a per-source basis. This SDSS-JAVELIN subsample is broadly similar to the published SDSS sample for the H and CIV lags, but prefers a markedly for lower scatter in the MgII relationship. For the MgII lags, these selection criteria drastically reduce the scatter of the lags about their best-fit relationship, and marginally decrease its preferred slope (see Figure 12(a) for a comparison).
5 OzDES Reverberation Mapping Results
In this section we summarise the reverberation mapping results across past OzDES papers, including the reverberation lags and the resulting estimated SMBH masses.
5.1 Lag measurements
In total, the OzDES reverberation mapping program has produced high quality AGN lags, consisting of H lags (OzDES-Malik_2023), MgII lags (OzDES-Yu_2021; OzDES-Yu_2023) and CIV lags (OzDES-Penton_2025). For each of these sources, we calculate the mass using equation 1 and accretion rate using equation 4, where the monochromatic luminosity uses bolometric corrections of Runnoe_2012 (Table 2 in their paper) and we use the virial factor of from SDSS-Shen_2023, as well as their dex inherent scatter. Rest-frame lags range from 333 being days, here and elsewhere in the paper. to , sitting within the range of lags recovered by other works. We find that the distribution of lags over redshift is reasonably consistent with existing reverberation mapping results from the literature (Figure 5, upper left). We note that the distribution of lags changes abruptly as one moves between lines, most notably between the MgII and CIV lines, indicating different lags, and thus radii, for different ions. This is quantified in Section 7. These RM results, as well as the redshifts and luminosities for the relevant sources, are listed in Table LABEL:tab:Results_All_Hbeta for our H sources, Table LABEL:tab:Results_All_MgII for MgII sources, and Table LABEL:tab:Results_All_CIV for CIV sources.
5.2 Black hole masses from direct RM measurements
In total, we present a compilation of the SMBH masses from OzDES reverberation mapping, ranging from to . The relevant measurements are listed, along with estimates of the corresponding accretion rates, in Tables LABEL:tab:Results_All_Hbeta, LABEL:tab:Results_All_MgII and LABEL:tab:Results_All_CIV. Accretion rates are calculated per equation 4, correcting luminosities to the using the bolometric corrections of Runnoe_2012. The calculation uses full error propagation by varying the lag, luminosity and line width within their respective measurement uncertainties. Results are reported as the media, with uncertainties drawn from the 16th and 84th percentiles.
We broadly see an increase in mass with redshift (Figure 5 lower left), which is expected, as the AGN at higher redshifts are selected to be more luminous (see, for example, Figure 1 in OzDES-Malik_2024), and this selection effect will dominate any evolutionary effects. We similarly see the expected general increase of more luminous AGN being more massive, but are unable to infer a clear power-law relationship (Figure 5 lower right).
It should be noted also that the estimated accretion rates for high redshift CIV sources can be extremely high (Figure 5 upper right), with most direct RM samples sitting above the threshold used by Du_2015 as the division between low and high accretion sources. In Section 6.1.3 we note an apparent flattening of the CIV curve at high luminosity, and, given that SEAMBH associate high accretion rates with low lags (HBETA_U_2022), this offers a possible physical explanation. Within the OzDES sample, there is a weak trend for more luminous, higher-redshift sources to accrete more strongly (black data points, Figure 5 upper right).
6 Comparison of New Results With Past Surveys & Tuning of Relationship
Here we examine the relationship using the full array of lags available at the close of this latest generation RM surveys. In addition to lags from OzDES and the final SDSS release of SDSS-Shen_2023, we use a wide array of past H, MgII and CIV lags as outlined in Section 3.3. In Section 6.1, we use a Bayesian approach to constrain the parameters of the relationship for all three reverberating emission lines, and to identify where the results of different surveys are in tension. For each line, we then select a ‘primary’ set of data for constraining an relationship that is most representative of the high redshift results of OzDES and SDSS, with the aim of demonstrating that such relations, marginalised over the footprint of such industrial scale surveys, are now within reach.
We investigate in Section 6.2 the possibility of using line width or accretion rate to reduce scatter in the single epoch model, and finally we use then use the relationship as a probe of the relative sizes of the H, MgII and CIV emission regions of the BLR in Section 7
6.1 Comparison of Parameter Constraints
In this section examine what constraints can be places on the relationship for all lines using OzDES data alone, and also using OzDES data in combination with data from other sources. Though OzDES and SDSS have similar physical target populations, the same can not be said for the wealth of literature sources. Further still, different surveys have different lag measurement methods and quality cut criteria, meaning we cannot assume them all to trace the same population in the plane. We resolve some of this issue by using the SDSS data products that most closely match the published OzDES population (see Section 4.2.4), but for other surveys we take a high level approach of only allowing data sets to be used in combination if there is some reasonable combination of parameters that satisfy all surveys in the grouping simultaneously. The rationale here is that every survey will have complex multi-layered selection effects, but we are here concerned only if these impose enough of a difference in the plane that they are significantly in tension with one another. A more complex treatment would require modelling of every survey, their selection effects and pipelines and the AGN population, while the approach taken here is a natural extension of our comparison of the surveys in the plane. In Table 2 we list the parameters for our chosen ‘primary’ data sets, while the following sub-sections elaborate on alternate relationships with different combinations of literature data and quality cuts.
For consistency, we refit the relationship with the same model for all data, while making use of the published lags and monochromatic luminosities from previous papers.444OzDES luminosities differ slightly from previous papers due to recalibration. We model the relationship as being linear in log-log space, with slope , offset , and an intrinsic scatter such that the a given luminosity predicts a lag by:
Here, represents the choice of luminosity units, typically or erg/s, and is chosen to minimise the covariance between offset and slope in the posterior distribution. varies with the choice of time units for the lag, and we here follow the convention of all lags being in days.
The observed RM lag, , has an additional scatter from observational uncertainty. We describe this uncertainty in log-space, i.e. . Under this modelling, the resulting likelihood for fit parameters is:
We approximate the measurement uncertainty, , as an asymmetric ‘split Gaussian’ with uncertainties found from the positive and negative uncertainties of the lag measurement. I.e. if a lag measurement is constrained within upper and lower bounds and , then the log-space measurement uncertainty is:
For similar consistency reasons, we use the same procedure to re-fit the relation of literature sources and combinations thereof. We treat all published lags and their uncertainties as describing the cumulative summary statistics of the lag posterior distribution (i.e. median and 16th, 84th percentiles).
All fits are performed using the Bayesian analysis tool NumPyro (numpyro), specifically its No U-Turn (NUTS) sampler, a Hamiltonian Monte Carlo (HMC) method that is well suited to unimodal distributions. We note that this fitting is relatively insensitive to the choice of sampler, as it is unimodal and produces approximately Gaussian contours (e.g. Figure 12(a)).
| Line | Luminosity Unit | Slope () | Offset () | Scatter () | Slope - Offset Fit Correlation () |
| H | |||||
| MgII | |||||
| CIV | |||||
6.1.1 H Relationship
OzDES is designed primarily for higher redshifts, and so captures only a small number of sources in the H redshift range. Coupled with the high rejection rate, this yields only high quality lag recoveries for the OzDES H relationship (OzDES-Malik_2023). In past OzDES analyses, the limited number of sources has been bolstered by lags from the literature, namely the 5-year SDSS lags (SDSS-Grier_2017), a wide range of low redshift lags collated by HBETA_Bentz_2009 as well as later additions to this ‘Bentz Collection’ (HBETA_Bentz_2013; HBETA_Bentz_2014; HBETA_Bentz_2016a; HBETA_Bentz_2016b; HBETA_Bentz_2023), lags from the SEAMBH project, lags from the LAMP collaboration (HBETA_U_2022), and a small number of miscellaneous literature lags (HBETA_Li_2021; HBETA_Lu_2016; HBETA_Zhang_2019; HBETA_Rakshit_2019). We make use of these same sources, in addition to the final SDSS lags of SDSS-Shen_2023, but re-fit each data set individually to more fairly measure tensions between them. Considering surveys to be in tension if they differ by more than , the groups of surveys with mutually consistent constraints are listed in Table 3. In total, OzDES sources and SDSS sources are used.
Figure 12(a) (a) shows relationship parameters for data from OzDES, SDSS, the Bentz Collection, the SEAMBH Survey, the LAMP survey, and a number of other miscellaneous sources. Shown also are the contours for SDSS lags as recovered from JAVELIN instead of PyROA, and selected according to the criteria outlined in Section 4.2.4. The constraining power is highest for surveys with a high source-count, like SDSS & SEAMBH, or a wide range of luminosities as in the Bentz Collection.
Due to the low sample count, the OzDES and miscellaneous sources are poorly constrained and are consistent with all other surveys. The largest tension exists between the slope and offset of the Bentz Collection and SEAMBH data, with this tension being the deciding factor about which surveys are mutually consistent. The SEAMBH data are known to produce consistently lower lags from their highly accreting AGN, and this is borne out by the lower values for and in collections including SEAMBH sources.
The SDSS results are consistent with these SEAMBH values in slope and scatter, but show systematically higher scatter than any other survey with strong constraints. This lower offset and higher scatter puts the SDSS data in tension with the low redshift Bentz Collection. Limiting the SDSS lags to only those that pass our cuts in Section 4.2.4 (the SDSS-JAVELIN contour), the constraints become weak enough to overcome this inconsistency, but the preferred values for , and do not significantly change. The LAMP survey lags sit the lowest of any data set, and their comparatively narrow range of luminosities mean they do not strongly constrain the slope by themselves.
Figure 12(a) (b) shows the constraints on relationship parameters when combining all datasets that are not in tension with one another. Four such groups are possible:
1.
OzDES combined with the Bentz Collection, miscellaneous sources and the sub-sampled SDSS data,
2.
As for 1, but with SEAMBH lags in place of Bentz,
3.
As for 2 but using the full SDSS sample,
4.
OzDES data combined with miscellaneous sources, the SEAMBH lags and the lags from the LAMP survey.
Where we use a ‘primary’ H grouping elsewhere in the paper, it refers to the first of these groupings. We make this choice of grouping as it combines the data of OzDES and SDSS with a broad set of literature lags, while excluding the lags of the SEAMBH survey that are, by the survey’s design, deliberately intended to be atypical of the broader AGN population.
| Source |
OzDES |
SDSS |
SDSS-JAV |
LAMP |
Bentz |
SEAMBH |
Misc |
| OzDES | - | T | T | T | T | T | T |
| SDSS | T | - | - | F | F | T | T |
| SDSS-JAV | T | - | - | F | T | T | T |
| LAMP | T | F | F | - | F | T | T |
| Bentz | T | F | T | F | - | F | T |
| SEAMBH | T | T | T | T | F | - | T |
| Misc | T | T | T | T | T | T | - |
6.1.2 MgII Relationship
Until this recent generation of high redshift surveys, the bulk of reverberation mapping lags have been measured using H, with a far more sparse set of literature lags in MgII. Rather than dividing these few sources by survey, we compare the OzDES and SDSS results to the convenient collection by MgII_Zajacek_2020, consisting of lags from MgII_Lira_2018, MgII_Czerny_2019 and MgII_Metzroth_2006, and a recent lag from MgII_Zajacek_2021. We exclude the lags from MgII_Shen_2016 and MgII_Shen_2019, as they are already included in the SDSS results sample. From the OzDES and SDSS data releases we use sources each.
Comparing relationship parameters for SDSS and OzDES, we find that they broadly agree in terms of the slope and offset of the fit, but that SDSS lags exhibit significantly higher scatter by roughly dex. The low number of literature sources in the Zajacek Collection means that its fit is poorly constrained, and so is not in tension with OzDES or SDSS. Restricting the SDSS results to the SDSS-JAV sample drastically reduces the number and constraining power, but also decreases the scatter significantly. This decrease in scatter removes the tension with OzDES, allowing these two data sets to be combined.
This yields two mutually consistent groupings, making use of the full SDSS data set and excluding OzDES or including OzDES but restricting the SDSS data. Owing to the complementary correlations of the data sets, both choices have near identical results for the main properties of the relation, but with a slightly higher scatter for the SDSS-based results (Figure 12(a) (c)). We proceed with the OzDES + SDSS-JAVELIN + Zajacek fit as our primary MgII relation as it is derived from the most diverse set of sources while still showing very low scatter about the mean relationship, suggestive of a lower false positive rate.
6.1.3 Relationship
Like MgII, there are only a small number of CIV lags outside of large-scale high redshift surveys. For these literature lags, we draw sources from the collation by CIV_Kaspi_2021, including low redshift AGN in which the CIV line is observed in the observer-frame ultra-violet range from CIV_Peterson_2005; CIV_Rosa_2015 and MgII_Metzroth_2006, as well as high redshift sources from MgII_Lira_2018 and high redshift sources new to the Kaspi paper itself. As with the MgII literature collation, we exclude the Kaspi sources from CIV_Grier_2019 and OzDES-Hoormann_2019 as they are already included in the SDSS and OzDES data.555Interestingly, only one of the two sources from OzDES-Hoormann_2019 makes it into the final OzDES sample. In that paper we used the down-weighting of seasonal gaps method employed by SDSS (equation 5), however when that is removed, the confidence in those sources drop and only one of them passes (at bronze level) the simulation-based criteria of OzDES-Penton_2021. We use high quality lags from OzDES and from SDSS.
Unlike the H and MgII relationship fits, it is much harder to find consistent agreement between different data sources. The literature lags sit systematically lower than the trends set by either OzDES or SDSS, and there is consistently greater scatter than exhibited for either H or MgII. Some of this discrepancy can be attributed to selection effects (see below); some may be due to CIV being contaminated by a non-virial wind component (Denney_2012). We attempt to account for selection effects in this section, but leave more detailed study of contamination for future work.
The high redshift surveys of SDSS and OzDES are highly consistent with one another except for the markedly higher scatter of the SDSS data, much as is seen for H and MgII. The OzDES-like SDSS-JAV sample has lower scatter, removing the tension with the OzDES sample, though partially through a loss of constraining power from the reduction in source count.
The literature lags sit noticeably lower than either SDSS or OzDES lags, even at high redshift. Though it is not statistically objectionable to combine both high redshift surveys with either the Kaspi Collection’s high redshift lags or its low redshift anchor, this gap means that it is not reasonable to do both at once. This allows for five possible groupings if taking the standard approach to the fitting of the relationship:
1.
The high redshift literature lags with the combined OzDES and SDSS-JAV sample,
2.
As above but with the low redshift anchor literature sources,
3.
As per 1. but with the full SDSS and no OzDES sources,
4.
As above but with the low redshift anchor literature sources,
5.
The low and high redshift sources of the Kaspi Collection with no new sources.
Though still exhibiting a high scatter, the above groupings that fit for the entire luminosity range (groups 1, 3 and 5) give slopes roughly consistent with those of H and MgII, while the flattened branch of the high-luminosity only groupings (2 and 4) have flat relationships between reverberation lag and AGN luminosity. This apparent disagreement between sources means we are unable to choose a primary data set for fitting the CIV relationship.
The tension between the slope of the high and low redshift is manifest in an apparent ‘levelling out’ of the relation at high redshift (Figure 6, bottom panel). This is easiest to see when plotting the residuals against the ‘high-’ fit (our ‘grouping 1)’). As shown in Figure 12, high luminosity / redshift sources give unmistakably and consistently higher lags than the lower luminosity / redshift data.
This nonlinear relationship for CIV may have a physical basis, such as distant AGN having higher accretion rates, which have been associated with lower lags (e.g. Du_2015; HBETA_Hu_2021). A simpler explanation lies in describing this as a selection effect: more luminous AGN have longer rest-frame lags, stretched further at high redshift by time dilation, and are therefore more likely to have lags that exceed the length of the survey. This induces a selection effect in which high redshift / luminosity sources are biased downwards from the true population distribution as the finite survey length imposes a ceiling on what observer-frame lags can be recovered. CIV sources are often the most distant, and consequently the most luminous and associated with the longest observer-frame lags, and so their observations are the most impacted by this effect.
Coupled with recovery rates and quality cuts, the exact effects of this under-sampling of high lag sources is difficult to characterise, as it interacts with the entire observational and analysis pipeline. A complete treatment would require modelling the underlying source population, survey selection, and the entire analysis chain including quality cuts. Though not intractable, this is an involved task, particularly when trying to unify multiple surveys and studies.
OzDES-Penton_2025 estimates that, for the -year baseline of OzDES, a lag-recovery selection effect becomes significant for CIV sources at observer-frame lags of and continues until lags above are almost impossible to recover. We make the coarse approximation of treating this limit as a hard cutoff, altering our Bayesian model so that that OzDES cannot detect lags above this limit. This is tantamount to altering the normal distribution of equation 7 to be a truncated normal distribution with a cutoff at . To a rough first order approximation, we scale this limit for each survey proportional to the length of their spectroscopic campaign as compared to OzDES, giving nominal lag ‘cutoffs’ in Table 4.
| Survey | Highest Observer Frame Lag (days) | Spectroscopic Baseline (yrs) | Nominal Lag Cutoff (days) |
| OzDES | |||
| SDSS-JAVELIN | |||
| Kaspi High-Z |
This model significantly relaxes the constraints from each survey, introducing the freedom for an un-observed fraction of sources to exist above the lag cutoff. This relaxes the tension between the surveys, creating a single parameter-space region of overlap between all four data sources (Figure 13, top panel). The high luminosity literature sources, drawn from the high- Kaspi sample, are dominated by this observational window effect, loosening their constraining power such that this fit matches the constraints of the more generic fit achieved when excluding the high- literature lags all together (Figure 13, bottom panel). It is this fit, with all CIV data accounted for, that we use as our primary fit for single-epoch mass estimates in Section 8.
6.2 Accretion Rate & Line Width as Predictors of RM Lag
It is matter falling on to the accretion disk circling the SMBH that drives AGN activity, and there is evidence of a separation in behaviour between low and high accretion AGN. Highly accreting ‘super-Eddington’ AGN are associated with lower lags than their ‘sub-Eddington’ counterparts at the same luminosity (Du_2015). It is also evident that high accretion sources tend to sit below our best fit relation (e.g. Figure 12), while low accretion sources sit above. We are then motivated to see if observations other than luminosity are predictors of AGN lag as a means to reduce the scatter of our single epoch methods, improve the predictive constraints of our single epoch methods and as a probe of the underlying AGN behaviour.
To examine this, we examine a new functional form for the single epoch lag prediction in which we add emission line width as a variable in the relation, i.e. an relation:
where , the slope of log-lag against log velocity dispersion, represents the predictive power of line-width independent of luminosity.
Our estimates for mass and accretion rate (Equation 1 and Equation 4, respectively) both follow power laws, making them linear for the logarithmic space that we fit our single epoch relations in. In this way, fitting for line width and luminosity also covers the fitting for accretion rate, as it depends entirely on these two other measurements.
Performing this fit for all lines, using our primary data groupings from Section 6.1, we find no significant improvement in predictive power (Table 5), and constraints on the lag-velocity dependence are broadly consistent with zero, i.e. no dependence of lag on line width. For MgII we see a slight preference towards , but this is only weakly constrained and there is still no marked decrease in scatter about the model.
| Line Type | for | for | |
| H | |||
| MgII | |||
| CIV |
The apparent trend between accretion rate and the residuals about our best fit relationships is a projection of the luminosity dependence onto the log-accretion rate axis, rather than an independent axis of variability. For a fixed luminosity, higher than average lags will give higher masses, which will give lower accretion rates, producing an apparent anti-correlation between the two.
There is the possibility that the lack of predictive power comes from a strong correlation between luminosity and velocity dispersion in our sample: if we only probe a narrow band of space, the two parameters become degenerate in our fitting. We can rule this out by examining the correlation between luminosity and velocity (Figure 14). We find that this correlation is weak, and is not consistent between lines. Table 6 notes an extremely shallow scaling index between luminosity and velocity dispersion for all lines, indicating that our sampling of the two parameters is sufficiently uncorrelated.
We note that this lack of predictive power in velocity is in contrast to the findings of the SEAMBH collaboration, (HBETA_Du_2016; HBETA_Du_2018; HBETA_Hu_2021) who find that higher accretion rates (derived from higher velocities) is associated with lower lags. Our results here do not necessarily contradict their findings, this analysis is for our primary data group which does not include their data specifically because their highly accreting sources produce statistically lower H lags (see Section 6.1.1), and we use a different functional form for the accretion dependence in place of their use of a critical accretion rate threshold. Our fit only demonstrates that line width / accretion rate does not serve to explain the residuals about our best fits. It is also worth noting that different data sources have non-uniform approaches to measuring these line widths, and unlike the lag parameter we make no attempt to account for these differences in analysis.
| Line Type | Log-Vel / Log-Lum Slope | Scatter (Dex) |
| H | ||
| MgII | ||
| CIV | ||
| All Sources |
7 Probing BLR Stratification by Comparisons of Bolometric Relations
In cases where there are lags for multiple lines from a single source (e.g. MgII_Metzroth_2006; MgII_Lira_2018), it has been observed that different RM lines produce different lags, indicating a stratification in the structure of the BLR, with different lines emitted from distinct regions. By comparing these lags, we can probe the relative size and overall geometry of the broad line region. Though a small subset of the OzDES target AGN do exhibit multiple lines in the optical range, these data are of insufficient number and quality to meaningfully measure the BLR stratification. Here we instead leverage the full set of available RM results by making use of the relationships themselves (as fit in Section 6.1) as a probe of the BLR geometry.
For each line, in Section 6.1 we derived relationships relative to the monochromatic luminosity evaluated at a line-specific wavelength conveniently similar to that of the reverberating line itself. Though each line’s relationship refers to a different luminosity, simple bolometric corrections allow us to reposition these relationships onto the same luminosity axis. Doing so allows us to compare the expected lags, and so typical radius of emission, for the different ions that produce reverberating emission lines. Here we use a variation of the relationship fitting in Section 6.1 to examine implications of the radii of the H, MgII and CIV emission regions.
In this analysis, we assume a fixed proportionality between monochromatic and bolometric luminosity:
We use the bolometric corrections of Runnoe_2012, namely that , , and . We use the correction for for CIV, though note that the luminosity is defined at . As we take the logarithm of these values, small variations should have little impact. The resulting bolometric luminosities span the range .
These corrections result in a horizontal shift of the relations that were derived in Section 6.1 and presented in Table 2. We note that this is the simplest possible bolometric model, applying only a simple translation of the relationship, while more expressive alternatives exist that describe the bolometric correction as a luminosity dependent power law would also impact the relationship’s slope. We note that it is a coarse assumption to treat these corrections as being equally applicable across all redshift ranges, and stress that this is a simplified first pass of this stratification measurement.
By correcting each source to its bolometric luminosity, we can repeat the fitting procedure of Section 6.1 and more meaningfully compare the relations for H, MgII and CIV. For all lines, we use the same data sets as the final fits in Table 2, including our treatment of selection effects for CIV as discussed in Section 6.1.3.
Performing this fit with luminosity units of (Figure 19(a) (b)), we find that the MgII and H relations are very similar, though the MgII lags sit at a slightly higher offset and prefer a slightly shallower slope. By contrast, the fit for CIV sits lower, but with a slope and scatter that is broadly similar to H. This is in keeping with prior studies of sources with multiple lines visible simultaneously, which found that H and MgII were produced by roughly co-spatial regions of the BLR (i.e. yielded similar lags for the same source, for an example see MgII_Shen_2019), while CIV lags tend to be smaller by a factor of , though this ratio is poorly constrained. (MgII_Lira_2018; Kaspi_2007).
We can constrain these relative radii for the different emission lines by adopting an adjusted relationship with dimensionless scaling factors ‘’ for MgII and CIV (equation 11):
Here, units for lag and luminosity are as in equation 6. This allows each line’s relation to have a different vertical offset, equivalent to the assumption that the ratio of these lags is the same across different AGN. 666A further model was trialled in which each line was free to fit its own inherent scatter, but this did not significantly change the resulting parameter constraints for slope or scale factors compared to when a single intrinsic scatter is fit.
By fitting this combined model we arrive at a bolometric relationship with a slope of . The constraints on this model give two interesting results:
1.
Constraints on suggest MgII produces longer lags than H, indicating that it may be emitted from a radius larger than H by a factor of . OzDES_Malik_2024 who did a stacked RM analysis (see their Table 1), in which they found that the average MgII lags in the OzDES sample were significantly longer than those for H at comparable luminosities.
2.
We find that the CIV radius is consistent with begin co-spatial with the H radius, though it prefers to be smaller by a factor of
The constraints on the relation as well as the relative scaling factors are summarised in Table 7, with a sketch of the implications for the BLR geometry shown in Figure 15. These findings are in line with those of SDSS-Shen_2023, who similarly found MgII lags to be systematically longer than those for H and CIV lags to be systematically shorter.
| Data | Slope () | Offset () | Scatter () | Slope - Offset Fit Correlation () | ||
| H | - | - | ||||
| H + MgII | - | |||||
| H + CIV | - | |||||
| H + MgII + CIV | ||||||
It is worth noting that the relative radius of H and MgII are strongly dependent on the assumptions of AGN spectral properties. Bolometric corrections are entirely degenerate with relative radius parameter , and the corrections of Runnoe_2012 are derived from only low redshift AGN. There is also considerable diversity amongst bolometric corrections from different sources (e.g. Runnoe_2012; Netzer_2019; Richards_2006), though there is broad agreement on the relative corrections between H and MgII, i.e. that .
8 Black Hole Masses From Single-Epoch Measurements (New Results)
We apply the relationships derived in Section 6.1 and summarised in Table 2, to derive black hole masses for OzDES AGN that do not have time-series spectra. We use the luminosity of the AGN to estimate their lag, and combine that with a measure of the line width at a single epoch to estimate their mass. The results are subject to a set of quality cuts to avoid contamination of poor quality spectra. We require that all single-epoch sources:
1.
Have a clearly visible line profile by visual inspection,
2.
Have a mass that excludes a nonphysical zero/negative mass at a level
After these cuts, we derive the masses of an additional OzDES AGN using these single-epoch estimates, consisting of H sources, MgII sources and CIV sources. This single epoch sample results in masses ranging from to . For the CIV sources, we make use of our ‘selection effect adjusted’ relationship (see Section 6.1 for details).
As with the RM sources in Section 5.2, we also use equation 4 to estimate the accretion rates of these sources, using the bolometric corrections of Runnoe_2012 to convert monochromatic luminosities to the required units. All single epoch sources, including their masses, accretion rates, single epoch radii and measured velocity dispersions, are listed in Table LABEL:tab:single_epoch. These masses are included, along with the RM and literature measurements, in Figure 5.
Masses are again calculated by varying all measurements within their uncertainties, with added variation for the single epoch lag within the uncertainties of our parameters. As with all other calculations, values are quoted as the distribution median, and uncertainties quoted using from 16th and 84th percentiles of the marginalised posterior distribution for each parameter. To avoid extrapolating relationships too far, monochromatic log-luminosities must be for H sources, for MgII sources and for CIV sources. The luminosity limit is somewhat arbitrary, so we include the single MgII source outside this range in Table LABEL:tab:single_epoch, marked with a †.
We find that the ceiling on the AGN SMBH masses has lowered towards more recent times, with no SMBH above nearer than in our results. If we limit ourselves to only the DES single-epoch data of Table LABEL:tab:single_epoch we acquire a single homogeneous data set that avoids the observation baseline dependent effects discussed in Section 6.1.3. Even for this data set, we still see a significant evolution of the mass-distribution of sources as a function of redshift (Figure 19), with nearby sources showing a consistent shift towards lower masses. As OzDES is a magnitude limited survey, some of the lack of distant low-mass observations can be explained by low completeness, dim objects being harder to observe at greater distances, but we can still confirm that the high-mass tail of the mass-density distribution, which should be subject only to count-based statistical effects, also decreases for nearby times to a degree that cannot be explained by statistical uncertainty or by a decrease in the number of available sources (see Figure 20). This is not a shocking result, as it aligns with the well established trend of ‘cosmic downsizing’ by which massive SMBH become less active towards more recent cosmic history (Barger_2005_downsizing; Vestergaard_2009_downsizing; Kelly_2010_downsizing; Fanidakis_2011_downsizing).
9 Discussion and Conclusions
This work marks the close of the OzDES reverberation mapping program, and, in conjunction with the final RM data release of the Sloan Digital Sky Survey by SDSS-Shen_2023, the end of the first generation of industrial scale high redshift reverberation mapping. OzDES has measured lags for AGN, including , and measurements from the H, MgII and CIV reverberating lines, respectively, and we have in this work presented an additional SMBH masses from single epoch methods. These SMBH masses represent a significant contribution to the number of high redshift black hole mass measurements. The SMBHs range in mass from to , putting our most massive measurement at the higher range of presently observed SMBHs.
Supplementing OzDES data with existing lags from the literature that are consistent OzDES’s high-redshift sources, we provide significantly improved constraints on the radius-luminosity relationship for sub-Eddington H and MgII lags. For all lines, we do not find a strong impact from accretion rate or emission line broadening as a single epoch lag predictor. Applying our MgII relationship to ULAS J1120+0641, a particularly distant SMBH observed by JWST at a redshift of and a luminosity of and MgII line FWHM of (Banados_2017), we recover a mass of , slightly lower, but still consistent with, Banados_2017’s estimate of .
Our single epoch relations for H and MgII have a significantly smaller scatter than those of the final SDSS release, dex against their dex. As these relations are tuned over a large redshift range, they offer a broadly applicable tool for estimating SMBH masses to higher redshifts. We also offer the best constrained CIV fit at high redshift, but we recommend some caution in extrapolating this relation beyond the luminosity range of the available RM-data (i.e. ).
Using our new relations, we make single epoch mass estimates for AGN up to redshift . In these estimates we are able to observe a dearth of massive AGN, with the upper range of AGN mass trending downwards at recent cosmic times. We note that measurements we present are subject to a selection function that depends on interaction between the physical properties of the AGN, the varied analysis pipelines of different surveys and the quality cuts that we apply here. These effects are particularly strong at high redshift where lags abut our upper observational limit of observer-frame lags. We make no attempt to model these selection effects, and so defer any interrogation of the black hole population properties to future work.
The similarity of the MgII and H relationship in slope and scatter, even with a statistically significant sample such as we now have access to, supports our understanding of their emission regions as being stratified. Of interest is the systematically larger MgII lags for comparable bolometric luminosity. Though simple arguments of ionisation energy suggest that the higher energy MgII should precipitate out closer in than H, some photoionisation models (e.g. Guo_2020_Photoionization) do support MgII being farther out from the SMBH engine. We note this with the caveat that our analysis is based on the simplest possible modelling of the spectral-energy distribution of AGN. Conversely, our fits indicate that CIV may be emitted from a radius much closer to that of H than found from simultaneous RM of H and CIV (MgII_Lira_2018; Kaspi_2007), though a slightly smaller radius is statistically preferred. The challenges in fitting a single relationship for this line make it difficult to constrain this with any certainty.
The strong windowing effects of the CIV trend highlights the need for further study of this emission line,777For example, to investigate whether the CIV line is affected by outflows, which could contaminate the line’s velocity dispersion measurement. either through intensive study of low redshift sources or by extending the luminosity envelope of CIV lag recoveries. Nearby, low luminosity sources can be observed in the observer-frame ultraviolet range, (e.g. MgII_Metzroth_2006; MgII_Lira_2018), and extending such studies to higher redshift or luminosities would help to clarify the origins of the apparent flattening of the CIV curve. For the same reason, there is also motivation to use infrared spectroscopy to observe the high redshift counterparts to the H and MgII sample to determine if these systems are homogeneous or show evolution over time. Extending these samples to more distant, more luminous sources will allow us to determine whether this ‘levelling out’ of the relationship is physical, a statistical anomaly, or a selection effect. It will also let us probe the evolution of the geometry of such systems over time, and help delineate between accretion and luminosity based effects. In particular, follow up of sources in the low redshift domain, where the lower impact of time dilation means AGN variability is fast and lags are short, will allow direct comparison between CIV lags to H lags from and to MgII lags from . MgII_Metzroth_2006 and MgII_Lira_2018 already do this for some local AGN, but extending to more distant and luminous sources will help fill the low luminosity branch of the CIV relationship.
The major findings of this paper are twofold: firstly that we have enough data as of this generation in BLR RM to make meaningful scaling relationships for H, MgII and CIV out to high redshift, and that the large apparent disagreement between surveys can be resolved with even a simple accounting of selection effects and a consistent framework for screening false positives. In the resulting relations the inherent scatter now vastly outstrips statistical uncertainty, meaning the next generation of BLR RM may concern itself with the second order effects of sub-populations, multi-variable lag predictors (e.g. accretion rate), sub-populations and more granular statistical biases. The second finding, and one we make an effort to stress, is that the handling of these selection effects and false positives are not a fully settled issue.
Though BLR RM surveys are conceptually simple, surveys can adopt wildly divergent approaches in the particulars of their lag recovery. Different regimes for fitting lags and different criteria for quantifying the significance of these fits interact with initial survey target selection in a complex and multi-layered way. In this paper we offer relations that are a significant improvement over the existing standard, specifically for high redshift sources. However there remains work to be done in future to improve these constraints, even with existing RM measurements. There are two main avenues of interest for strengthening our RM-derived scaling relations in their current form: firstly, to re-fit lags for all sources across all surveys in a single framework, and secondly to properly model how this framework interacts with surveys to understand what selection functions arise in the luminosity / redshift / lag plane. We suggest that characterising these window functions should include a combination of simulation-based forward modelling in the style of (OzDES-Penton_2021) with modern fitters like LITMUS and PyROA do not suffer from JAVELIN’s numerical artefacts when encountering aliasing (see McDougall_2025_LITMUS). Of particular interest to the OzDES data would be a re-analysis that discards less false negatives for a larger final sample, and understanding how such cuts affects the apparent population distributions in addition to the overall false positive rate.
Contribution Statement
Project conception and coordination: TMD, CL, PM; Analysis, programming, calculations: HM with input from ZY, UM, AP; Writing: HM; Original figures: HM; Editing: TMD, CL, PM, ZY, RS, GL, BT, BP; Data generation and/or curation: all authors.
Acknowledgements
HGM, TMD, AP acknowledge support for early stages of this project from an Australian Research Council (ARC) Laureate Fellowship (project number FL180100168) and for later stages from the ARC Centre of Excellence for Gravitational Wave Discovery, OzGrav (CE230100016).
Funding for the DES Projects has been provided by the U.S. Department of Energy, the U.S. National Science Foundation, the Ministry of Science and Education of Spain, the Science and Technology Facilities Council of the United Kingdom, the Higher Education Funding Council for England, the National Center for Supercomputing Applications at the University of Illinois at Urbana-Champaign, the Kavli Institute of Cosmological Physics at the University of Chicago, the Center for Cosmology and Astro-Particle Physics at the Ohio State University, the Mitchell Institute for Fundamental Physics and Astronomy at Texas A&M University, Financiadora de Estudos e Projetos, Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro, Conselho Nacional de Desenvolvimento Científico e Tecnológico and the Ministério da Ciência, Tecnologia e Inovação, the Deutsche Forschungsgemeinschaft and the Collaborating Institutions in the Dark Energy Survey.
The Collaborating Institutions are Argonne National Laboratory, the University of California at Santa Cruz, the University of Cambridge, Centro de Investigaciones Energéticas, Medioambientales y Tecnológicas-Madrid, the University of Chicago, University College London, the DES-Brazil Consortium, the University of Edinburgh, the Eidgenössische Technische Hochschule (ETH) Zürich, Fermi National Accelerator Laboratory, the University of Illinois at Urbana-Champaign, the Institut de Ciències de l’Espai (IEEC/CSIC), the Institut de Física d’Altes Energies, Lawrence Berkeley National Laboratory, the Ludwig-Maximilians Universität München and the associated Excellence Cluster Universe, the University of Michigan, NSF NOIRLab, the University of Nottingham, The Ohio State University, the University of Pennsylvania, the University of Portsmouth, SLAC National Accelerator Laboratory, Stanford University, the University of Sussex, Texas A&M University, and the OzDES Membership Consortium.
Based in part on observations at NSF Cerro Tololo Inter-American Observatory at NSF NOIRLab (NOIRLab Prop. ID 2012B-0001; PI: J. Frieman), which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation.
The DES data management system is supported by the National Science Foundation under Grant Numbers AST-1138766 and AST-1536171. The DES participants from Spanish institutions are partially supported by MICINN under grants PID2021-123012, PID2021-128989 PID2022-141079, SEV-2016-0588, CEX2020-001058-M and CEX2020-001007-S, some of which include ERDF funds from the European Union. IFAE is partially funded by the CERCA program of the Generalitat de Catalunya.
We acknowledge support from the Brazilian Instituto Nacional de Ciência e Tecnologia (INCT) do e-Universo (CNPq grant 465376/2014-2).
This document was prepared by the DES Collaboration using the resources of the Fermi National Accelerator Laboratory (Fermilab), a U.S. Department of Energy, Office of Science, Office of High Energy Physics HEP User Facility. Fermilab is managed by Fermi Forward Discovery Group, LLC, acting under Contract No. 89243024CSC000002.
Based in part on data acquired at the Anglo-Australian Telescope, under program Ab/2013B/012]. We acknowledge the traditional owners of the land on which the AAT stands, the Gamilaroi people, and pay our respects to elders past and present.
Calculations were made using python (VanRossum_2009_python) and with the aid of numpy (harris_2020_numpy). Plots and figures were generated with the aid of matplotlib (Hunter_2007_matplotlib) and chainconsumer (Hinton_2016_chainconsumer).
We acknowledge and pay respect to the traditional owners of the land on which the University of Queensland and University of Southern Queensland are situated, upon whose unceded, sovereign, ancestral lands we work. We pay respects to their Ancestors and descendants, who continue cultural and spiritual connections to Country.
Based in part on data acquired at the Anglo-Australian Telescope, under program A/2013B/12. We acknowledge the traditional custodians of the land on which the AAT stands, the Gamilaraay people, and pay our respects to elders past and present.
Data Availability
With this paper we release a table of the OzDES RM light curves for the full sample, as well as tabulated forms of the RM and single epoch samples listed in Appendix A (Tables LABEL:tab:Results_All_Hbeta, LABEL:tab:Results_All_MgII, LABEL:tab:Results_All_CIV), and Appendix C (Table LABEL:tab:single_epoch). This is available at [link to be provided upon acceptance].
The raw images and catalogues from the whole DES survey are available at https://des.ncsa.illinois.edu/releases/dr2.
The spectra are publicly available from https://docs.datacentral.org.au/ozdes/overview/dr2/.
Affiliations
1School of Mathematics and Physics, University of Queensland, Brisbane, QLD 4072, Australia, 2Kavli Institute for Particle Astrophysics & Cosmology, P. O. Box 2450, Stanford University, Stanford, CA 94305, USA, 3Center for Cosmology and Astro-Particle Physics, The Ohio State University, Columbus, OH 43210, USA, 4Department of Astronomy, The Ohio State University, Columbus, OH 43210, USA, 5Centre for Gravitational Astrophysics, College of Science, The Australian National University, ACT 2601, Australia, 6The Research School of Astronomy and Astrophysics, Australian National University, ACT 2601, Australia, 7Sydney Institute for Astronomy, School of Physics, A28, The University of Sydney, NSW 2006, Australia, 8School of Mathematical & Physical Sciences, 12 Wally’s Walk, Macquarie University, Macquarie Park, NSW 2113, 9Fermi National Accelerator Laboratory, P. O. Box 500, Batavia, IL 60510, USA, 10Physik-Institut, University of Zúrich, Winterthurerstrasse 190, CH-8057 Zúrich, Switzerland, 11Departamento de Física Teórica, Centro de Astropartículas y Física de Altas Energías (CAPA), Universidad de Zaragoza, 50009 Zaragoza, Spain, 12Institute of Cosmology and Gravitation, University of Portsmouth, Portsmouth, PO1 3FX, UK, 13University Observatory, LMU Faculty of Physics, Scheinerstr. 1, 81679 Munich, Germany, 14Department of Physics & Astronomy, University College London, Gower Street, London, WC1E 6BT, UK, 15Instituto de Astrofisica de Canarias, E-38205 La Laguna, Tenerife, Spain, 16Laboratório Interinstitucional de e-Astronomia - LIneA, Av. Pastor Martin Luther King Jr, 126 Del Castilho, Nova América Offices, Torre 3000/sala 817, Brazil, 17Universidad de La Laguna, Dpto. Astrofísica, E-38206 La Laguna, Tenerife, Spain, 18INAF-Osservatorio Astronomico di Trieste, via G. B. Tiepolo 11, I-34143 Trieste, Italy, 19Korea Astronomy and Space Science Institute, 776 Daedeok-daero, Yuseong-gu, Daejeon 34055, South Korea, 20Institut de Física d’Altes Energies (IFAE), The Barcelona Institute of Science and Technology, Campus UAB, 08193 Bellaterra (Barcelona) Spain, 21Kapteyn Astronomical Institute, University of Groningen, Landleven 12 (Kapteynborg, 5419), 9747 AD Groningen, The Netherlands, 22Hamburger Sternwarte, Universität Hamburg, Gojenbergsweg 112, 21029 Hamburg, Germany, 23Centro de Investigaciones Energéticas, Medioambientales y Tecnológicas (CIEMAT), Madrid, Spain, 24California Institute of Technology, 1200 East California Blvd, MC 249-17, Pasadena, CA 91125, USA, 25Instituto de Física Teórica UAM/CSIC, Universidad Autónoma de Madrid, 28049 Madrid, Spain, 26Centre for Astrophysics & Supercomputing, Swinburne University of Technology, Victoria 3122, Australia, 27Santa Cruz Institute for Particle Physics, Santa Cruz, CA 95064, USA, 28Center for Astrophysics Harvard & Smithsonian, 60 Garden Street, Cambridge, MA 02138, USA, 29Lawrence Berkeley National Laboratory, 1 Cyclotron Road, Berkeley, CA 94720, USA, 30Lowell Observatory, 1400 Mars Hill Rd, Flagstaff, AZ 86001, USA, 31Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Dr., Pasadena, CA 91109, USA, 32Department of Physics and Astronomy, University of Pennsylvania, Philadelphia, PA 19104, USA, 33George P. and Cynthia Woods Mitchell Institute for Fundamental Physics and Astronomy, and Department of Physics and Astronomy, Texas A&M University, College Station, TX 77843, USA, 34Université Grenoble Alpes, CNRS, LPSC-IN2P3, 38000 Grenoble, France, 35Center for Astrophysical Surveys, National Center for Supercomputing Applications, 1205 West Clark St., Urbana, IL 61801, USA, 36Department of Astronomy, University of Illinois at Urbana-Champaign, 1002 W. Green Street, Urbana, IL 61801, USA, 37Institució Catalana de Recerca i Estudis Avançats, E-08010 Barcelona, Spain, 38Department of Astrophysical Sciences, Princeton University, Peyton Hall, Princeton, NJ 08544, USA, 39School of Mathematics and Physics, University of Surrey, Guildford, Surrey, GU2 7XH, UK, 40Observatório Nacional, Rua Gal. José Cristino 77, Rio de Janeiro, RJ - 20921-400, Brazil, 41Ruhr University Bochum, Faculty of Physics and Astronomy, Astronomical Institute, German Centre for Cosmological Lensing, 44780 Bochum, Germany, 42Physics Department, Lancaster University, Lancaster, LA1 4YB, UK, 43Computer Science and Mathematics Division, Oak Ridge National Laboratory, Oak Ridge, TN 37831, 44Department of Astronomy and Astrophysics, University of Chicago, Chicago, IL 60637, USA, 45Cerro Tololo Inter-American Observatory, NSF’s National Optical-Infrared Astronomy Research Laboratory, Casilla 603, La Serena, Chile, 46Berkeley Center for Cosmological Physics, Department of Physics, University of California, Berkeley, CA 94720, US
Appendix A Lag recovery methods
The main two lag recovery methods are the interpolated cross-correlation function and Javelin, which we describe here.
A.1 PyCCF
Introduced to the field of reverberation mapping by ICCF-Gaskell_1987, the Interpolated Cross-Correlation Function (ICCF) is a purely non-parametric lag recovery method that relies on the cross-correlation () of the two light curves as they are shifted against one another in the time domain. The procedure is to linearly interpolate one (or both) of the light curves between observations and measure the Pearson correlation,
where and represent the sets of photometric and spectroscopic amplitudes respectively, and angled brackets indicate an inner product, approximated by a sum.888, where the overline represents the mean. The physical lag is then taken to be the delay that maximises this correlation, while the uncertainties are estimated by ‘bootstrapping’: recovering the peak correlation-lag from many realisations generated by randomly sub-sampling the observations and then re-sampling these observations within their measurement uncertainties.
The ICCF method is numerically low cost and robust, and has seen wide use in RM, particularly in low redshift campaigns with high cadence measurements (e.g HBETA_Barth_2013; HBETA_Rakshit_2019; HBETA_Zhang_2019). Through simulations, it has been found to agree with more rigorous models like JAVELIN to within statistical bounds, though with higher reported uncertainties (OZDES-Yu_2019). The low numerical cost of the ICCF means that it can easily be used as a diagnostic tool for testing the reliability of other more involved methods, as the test statistic gives a general ‘goodness of fit’ at any lag, and the low cost of evaluation means the entire lag parameter space can be searched exhaustively. OzDES makes use of PyCCF (PyCCF-Mouyan_2018), a vectorised python-based implementation of ICCF. OzDES-Malik_2023 estimates a false positive rate / -value for each lag by shuffling the time ordering of measurements, creating mock observations of signals with the same stationary statistics as the observed light curve. These were then used to tune the quality cuts outlined in Section 4.2.1.
ICCF is robust and ‘light-weight’, but has shortfalls in its lack of precision and its inability to adequately characterise the vague constraints of light curves where large gaps exist between observations, such as the six month seasonal gaps of OzDES and SDSS (OzDES-Malik_2022). The quality of lag recovery degrades rapidly as measurement cadence or observational season length decrease, as the linear interpolation lends false confidence while failing to capture the stochastic variations that take place between measurements. By contrast, the SDSS team’s preferred lag recovery method, PyROA (pyroa-ferus_2021), interpolates the light curves by adopting a ‘rolling average’ of the observations within a window of variable width, giving conservatively broad constraints within seasonal gaps compared to PyCCF’s spuriously tight ones. Non-parametric methods like PyCCF and PyROA allow a certain degree of flexibility to account for their approximate modelling, and so naturally inherit a degree of robustness against outliers or under-estimated errors (OZDES-Yu_2019).
A.2 JAVELIN
Where non-parametric methods interpolate between observations with as few assumptions about the underlying light curves as possible, the second class of fitting methods leverages our understanding of the underlying stochasticity of AGN variability. Though AGN do not follow a consistent light curve shape like some astrophysical objects (e.g. supernovae, Guy_2005), they do exhibit consistent statistical properties in their signals, namely a set power spectrum for the variations in their brightness. Specifically, these variations are known to follow a ‘damped random walk’ (DRW) (Kelly_2009; Kozlowski_2010a; MacLeod_2010), with only small deviations at very long and short timescales of variation (Zu_2013). The DRW is an example of a Gaussian process, specifically a first order continuous auto-regressive process, yielding a red-noise like scaling at high frequencies. This Gaussian-process-like structure in the stochastic variations creates statistical correlations within and between light curves, allowing us to interpolate between measurements in a physically principled way. For the DRW specifically, the auto-correlation is encoded in a covariance matrix, whose elements are defined as:
where are the variabilities of the light curves that measurements and belong to, are their measurement uncertainties, are their measurement times after the spectroscopic measurements offset by the lag , is the timescale of variability in the damped random walk and is the Kronecker delta.
The predominant Gaussian process modelling program at time of writing is the python-based program JAVELIN (JAVELIN-Zu_2010), a DRW-based implementation of the method outlined by Press_Rybicki_1989. This method models the AGN light curves in a Bayesian way, constructing a unified light curve from the photometric and spectroscopic observations by treating the response as a shifted, scaled and smoothed copy of the driving continuum. This approach gives a closed form likelihood function:
where is the covariance matrix defined in equation 13 and are the light curve measurements (photometric and spectroscopic) after offsetting by the mean of their respective light curves and is the total number of observations across all light curves. This likelihood allows the signal parameters, including the lag, to be constrained in a Bayesian framework. This procedure uses all available information simultaneously to constrain the light curve behaviour between measurements, more accurately characterising the uncertainty arising from large seasonal gaps.
Equation 14 requires the inversion of a large matrix of rank . Though the near-diagonal shape of equation 13 ameliorates this cost somewhat, this is still a computationally expensive process. To reduce the number of evaluations, JAVELIN uses the MCMC approach using the python-based package emcee (emcee-Foreman_Mackey_2013), which uses the Affine-Invariant Ensemble Sampler (AffineInvariant-Goodman_2010) to direct the bulk of its samples to high likelihood regions of parameter space. The seasonal windowing function of equatorial surveys, and the resulting aliasing problem (see Section 2.2 for details) can give rise to multi-modal distributions along the lag axis of parameter-space (e.g. CIV_Grier_2019), which can result in poor efficiency and inaccurate posterior distributions from emcee. Gaussian-process-based modelling is also sensitive to outlier data-points or under-estimated measurement uncertainties, which can spuriously over-constrain the light curve and disrupt the lag recovery results.
Another Gaussian process based lag recovery program similar to JAVELIN is CREAM (Starkey_2015_CREAM), which broadly adopts the same methodology as JAVELIN but assumes different Gaussian process statistics and parameterises its Bayesian priors in terms of physical properties (e.g. mass, accretion rate etc). CREAM has seen use in earlier SDSS papers, (e.g. CIV_Grier_2019), and includes an ‘error scaling’ parameter in its generative model to relax the light curve constraints and side-step the impacts of under-estimated uncertainties. Outlier rejection, for both JAVELIN and CREAM, still relies on ‘by-eye’ identification and removal of suspicious measurements (e.g. OzDES-Yu_2023).
Appendix B OzDES Reverberation Mapping Results
| Source ID | Redshift | Monochromatic Luminosity | Rest-Frame Lag (Days) | Velocity Dispersion (km/s) | Mass () | Dimensionless Accretion Rate |
| 0.127 | ||||||
| 0.237 | ||||||
| 0.310 | ||||||
| 0.314 | ||||||
| 0.332 | ||||||
| 0.354 | ||||||
| 0.644 | ||||||
| 0.707 |
| Source ID | Redshift | Monochromatic Luminosity | Rest-Frame Lag (Days) | Velocity Dispersion (km/s) | Mass () | Dimensionless Accretion Rate |
| 0.840 | ||||||
| 0.890 | ||||||
| 1.000 | ||||||
| 1.070 | ||||||
| 1.150 | ||||||
| 1.240 | ||||||
| 1.370 | ||||||
| 1.420 | ||||||
| 1.430 | ||||||
| 1.440 | ||||||
| 1.480 | ||||||
| 1.490 | ||||||
| 1.490 | ||||||
| 1.530 | ||||||
| 1.530 | ||||||
| 1.620 | ||||||
| 1.640 | ||||||
| 1.640 | ||||||
| 1.640 | ||||||
| 1.650 | ||||||
| 1.660 | ||||||
| 1.680 | ||||||
| 1.750 | ||||||
| 1.790 | ||||||
| 1.860 |
| Source ID | Redshift | Monochromatic Luminosity | Rest-Frame Lag (Days) | Velocity Dispersion (km/s) | Mass () | Dimensionless Accretion Rate |
| 1.922 | ||||||
| 1.928 | ||||||
| 1.936 | ||||||
| 1.952 | ||||||
| 1.988 | ||||||
| 2.031 | ||||||
| 2.041 | ||||||
| 2.092 | ||||||
| 2.096 | ||||||
| 2.150 | ||||||
| 2.188 | ||||||
| 2.219 | ||||||
| 2.257 | ||||||
| 2.283 | ||||||
| 2.315 | ||||||
| 2.328 | ||||||
| 2.384 | ||||||
| 2.425 | ||||||
| 2.452 | ||||||
| 2.459 | - | - | - | |||
| 2.462 | ||||||
| 2.492 | ||||||
| 2.500 | ||||||
| 2.563 | ||||||
| 2.593 | ||||||
| 2.677 | ||||||
| 2.745 | ||||||
| 2.773 | - | - | - | |||
| 3.451 |
Appendix C OzDES Single Epoch Results
| Source ID | Redshift | Line | Monochromatic Luminosity | Single Epoch Lag (Days) | Velocity Dispersion (km/s) | Single Epoch Mass () | Dimensionless Accretion Rate |
| 0.040 | H | ||||||
| 0.180 | H | ||||||
| 0.220 | H | ||||||
| 0.250 | H | ||||||
| 0.310 | H | ||||||
| 0.430 | H | ||||||
| 0.490 | H | ||||||
| 0.510 | H | ||||||
| 0.540 | H | ||||||
| 0.550 | H | ||||||
| 0.570 | H | ||||||
| 0.590 | H | ||||||
| 0.610 | H | ||||||
| 0.620 | H | ||||||
| 0.640 | H | ||||||
| 0.650 | H | ||||||
| 0.660 | H | ||||||
| 0.670 | H | ||||||
| 0.670 | H | ||||||
| 0.710 | MgII | ||||||
| 0.720 | MgII | ||||||
| 0.720 | MgII | ||||||
| 0.730 | MgII | ||||||
| 0.730 | MgII | ||||||
| 0.730 | MgII | ||||||
| 0.760 | MgII | ||||||
| 0.770 | MgII | ||||||
| 0.790 | MgII | ||||||
| 0.790 | MgII | ||||||
| 0.800 | MgII | ||||||
| 0.810 | MgII | ||||||
| 0.820 | MgII | ||||||
| 0.830 | MgII | ||||||
| 0.840 | MgII | ||||||
| 0.860 | MgII | ||||||
| 0.860 | MgII | ||||||
| 0.870 | MgII | ||||||
| 0.880 | MgII | ||||||
| 0.890 | MgII | ||||||
| 0.900 | MgII | ||||||
| 0.900 | MgII | ||||||
| 0.910 | MgII | ||||||
| 0.910 | MgII | ||||||
| 0.920 | MgII | ||||||
| 0.920 | MgII | ||||||
| 0.930 | MgII | ||||||
| 0.930 | MgII | ||||||
| 0.940 | MgII | ||||||
| 0.940 | MgII | ||||||
| 0.940 | MgII | ||||||
| 0.940 | MgII | ||||||
| 0.950 | MgII | ||||||
| 0.950 | MgII | ||||||
| 0.950 | MgII | ||||||
| 0.960 | MgII | ||||||
| 0.960 | MgII | ||||||
| 0.960 | MgII | ||||||
| 0.960 | MgII | ||||||
| 0.960 | MgII | ||||||
| 0.960 | MgII | ||||||
| 0.970 | MgII | ||||||
| 0.970 | MgII | ||||||
| 0.970 | MgII | ||||||
| 0.970 | MgII | ||||||
| 0.980 | MgII | ||||||
| 0.980 | MgII | ||||||
| 0.980 | MgII | ||||||
| 0.980 | MgII | ||||||
| 0.980 | MgII | ||||||
| 0.990 | MgII | ||||||
| 0.990 | MgII | ||||||
| 1.010 | MgII | ||||||
| 1.010 | MgII | ||||||
| 1.020 | MgII | ||||||
| 1.050 | MgII | ||||||
| 1.070 | MgII | ||||||
| 1.080 | MgII | ||||||
| 1.090 | MgII | ||||||
| 1.090 | MgII | ||||||
| 1.090 | MgII | ||||||
| 1.090 | MgII | ||||||
| 1.090 | MgII | ||||||
| 1.090 | MgII | ||||||
| 1.110 | MgII | ||||||
| 1.110 | MgII | ||||||
| 1.120 | MgII | ||||||
| 1.130 | MgII | ||||||
| 1.140 | MgII | ||||||
| 1.160 | MgII | ||||||
| 1.160 | MgII | ||||||
| 1.160 | MgII | ||||||
| 1.170 | MgII | ||||||
| 1.170 | MgII | ||||||
| 1.180 | MgII | ||||||
| 1.190 | MgII | ||||||
| 1.190 | MgII | ||||||
| 1.200 | MgII | ||||||
| 1.200 | MgII | ||||||
| 1.200 | MgII | ||||||
| 1.210 | MgII | ||||||
| 1.220 | MgII | ||||||
| 1.220 | MgII | ||||||
| 1.230 | MgII | ||||||
| 1.230 | MgII | ||||||
| 1.230 | MgII | ||||||
| 1.240 | MgII | ||||||
| 1.240 | MgII | ||||||
| 1.250 | MgII | ||||||
| 1.250 | MgII | ||||||
| 1.250 | MgII | ||||||
| 1.250 | MgII | ||||||
| 1.260 | MgII | ||||||
| 1.270 | MgII | ||||||
| 1.280 | MgII | ||||||
| 1.280 | MgII | ||||||
| 1.290 | MgII | ||||||
| 1.290 | MgII | ||||||
| 1.300 | MgII | ||||||
| 1.300 | MgII | ||||||
| 1.320 | MgII | ||||||
| 1.320 | MgII | ||||||
| 1.330 | MgII | ||||||
| 1.340 | MgII | ||||||
| 1.340 | MgII | ||||||
| 1.350 | MgII | ||||||
| 1.350 | MgII | ||||||
| 1.350 | MgII | ||||||
| 1.360 | MgII | ||||||
| 1.360 | MgII | ||||||
| 1.370 | MgII | ||||||
| 1.370 | MgII | ||||||
| 1.370 | MgII | ||||||
| 1.370 | MgII | ||||||
| 1.380 | MgII | ||||||
| 1.380 | MgII | ||||||
| 1.390 | MgII | ||||||
| 1.390 | MgII | ||||||
| 1.390 | MgII | ||||||
| 1.390 | MgII | ||||||
| 1.390 | MgII | ||||||
| 1.400 | MgII | ||||||
| 1.400 | MgII | ||||||
| 1.420 | MgII | ||||||
| 1.420 | MgII | ||||||
| 1.420 | MgII | ||||||
| 1.420 | MgII | ||||||
| 1.430 | MgII | ||||||
| 1.430 | MgII | ||||||
| 1.430 | MgII | ||||||
| 1.440 | MgII | ||||||
| 1.450 | MgII | ||||||
| 1.450 | MgII | ||||||
| 1.450 | MgII | ||||||
| 1.460 | MgII | ||||||
| 1.460 | MgII | ||||||
| 1.460 | MgII | ||||||
| 1.470 | MgII | ||||||
| 1.470 | MgII | ||||||
| 1.470 | MgII | ||||||
| 1.470 | MgII | ||||||
| 1.480 | MgII | ||||||
| 1.490 | MgII | ||||||
| 1.490 | MgII | ||||||
| 1.490 | MgII | ||||||
| 1.490 | MgII | ||||||
| 1.490 | MgII | ||||||
| 1.500 | MgII | ||||||
| 1.500 | MgII | ||||||
| 1.530 | MgII | ||||||
| 1.530 | MgII | ||||||
| 1.530 | MgII | ||||||
| 1.540 | MgII | ||||||
| 1.550 | MgII | ||||||
| 1.550 | MgII | ||||||
| 1.560 | MgII | ||||||
| 1.560 | MgII | ||||||
| 1.570 | MgII | ||||||
| 1.580 | MgII | ||||||
| 1.590 | MgII | ||||||
| 1.590 | MgII | ||||||
| 1.600 | MgII | ||||||
| 1.600 | MgII | ||||||
| 1.610 | MgII | ||||||
| 1.610 | MgII | ||||||
| 1.620 | MgII | ||||||
| 1.620 | MgII | ||||||
| 1.630 | MgII | ||||||
| 1.640 | MgII | ||||||
| 1.650 | MgII | ||||||
| 1.650 | MgII | ||||||
| 1.680 | MgII | ||||||
| 1.680 | MgII | ||||||
| 1.680 | MgII | ||||||
| 1.680 | MgII | ||||||
| 1.690 | MgII | ||||||
| 1.690 | MgII | ||||||
| 1.690 | MgII | ||||||
| 1.700 | MgII | ||||||
| 1.700 | MgII | ||||||
| 1.710 | MgII | ||||||
| 1.720 | MgII | ||||||
| 1.730 | MgII | ||||||
| 1.730 | MgII | ||||||
| 1.750 | MgII | ||||||
| 1.760 | MgII | ||||||
| 1.800 | MgII | ||||||
| 1.800 | CIV | ||||||
| 1.830 | MgII | ||||||
| 1.830 | CIV | ||||||
| 1.830 | MgII | ||||||
| 1.830 | CIV | ||||||
| 1.830 | MgII | ||||||
| 1.830 | CIV | ||||||
| 1.850 | MgII | ||||||
| 1.850 | CIV | ||||||
| 1.860 | MgII | ||||||
| 1.870 | MgII | ||||||
| 1.870 | CIV | ||||||
| 1.880 | MgII | ||||||
| 1.890 | MgII | ||||||
| 1.890 | CIV | ||||||
| 1.900 | MgII | ||||||
| 1.900 | CIV | ||||||
| 1.910 | MgII | ||||||
| 1.920 | CIV | ||||||
| 1.920 | MgII | ||||||
| 1.930 | CIV | ||||||
| 1.970 | CIV | ||||||
| 1.990 | CIV | ||||||
| 2.030 | CIV | ||||||
| 2.060 | CIV | ||||||
| 2.070 | CIV | ||||||
| 2.080 | CIV | ||||||
| 2.100 | CIV | ||||||
| 2.110 | CIV | ||||||
| 2.200 | CIV | ||||||
| 2.220 | CIV | ||||||
| 2.230 | CIV | ||||||
| 2.240 | CIV | ||||||
| 2.250 | CIV | ||||||
| 2.250 | CIV | ||||||
| 2.260 | CIV | ||||||
| 2.290 | CIV | ||||||
| 2.390 | CIV | ||||||
| 2.440 | CIV | ||||||
| 2.480 | CIV |